OpenStudy (anonymous):

For 2x - 3y = -7, when x = 4, What is the value of y? Choose one answer. a. -1/3 b. 1/3 c. -5 d. 5

OpenStudy (anonymous):

just put value of x in your equation: 2*4-3y=-7 can you solve it?

OpenStudy (anonymous):

hm i think so :)

OpenStudy (anonymous):

give me your answer - will check it together :)

OpenStudy (anonymous):

hmmm i think it may be 5? but im probs wrong... :/

OpenStudy (anonymous):

heres a hint - If you substitute x = 4 into the equation you will get 2(4) - 3y = -7 or equivalently 8 - 3y = -7

OpenStudy (anonymous):

you are right! -3y=-7-8 y=-15/ (-3) y=5 :))

OpenStudy (anonymous):

Your answer of 5 is correct.

OpenStudy (anonymous):

omg! yay!!! :)

OpenStudy (anonymous):

option d is correct.... 8-3v=-7 3v=8+7 3v=15 v=5